# Asymptotic Stability of 3D Out-flowing Compressible Viscous Fluid under Non-Spherical Perturbation

### v1

@@ 2 Main Theorem
    We also denote the norms
    $\|F\|_{k}\vcentcolon=\bigg{(}\int_{\Omega}|D^{k}F|^{2}{\mathrm{d}}x\bigg{)}^{\frac{1}{2}},\qquad\|F\|\vcentcolon=\|F\|_{0},\qquad\|F\|_{\infty}\vcentcolon=\sup\limits_{x\in\Omega}|F|.$
3 A-Priori Estimates
Define the difference functions:
$\phi\vcentcolon=\rho-\tilde{\rho},\qquad\psi\vcentcolon=u-\tilde{u},\qquad(\phi_{0},\psi_{0})\vcentcolon=(\phi,\psi)|_{t=0},$
and denote $D_{t}\vcentcolon=\partial_{t}+u\cdot\nabla$ as the material derivative. Then (1.1) is rewritten as
$\mathcal{L}^{0}(\phi,\psi)\vcentcolon=$ $D_{t}\phi+\rho_{+}\textnormal{{div}}\psi=f^{0},$ (3.8a)
$\mathcal{L}^{i}(\phi,\psi)\vcentcolon=$ $\partial_{t}\psi^{i}-\frac{\mu}{\rho_{+}}\Delta\psi^{i}-\frac{\mu+\lambda}{\rho_{+}}\partial_{x_{i}}\textnormal{{div}}\psi+q_{+}\partial_{x_{i}}\phi=f^{i},$ (3.8b)
for $i=1,2,3$, where $q(\rho)\vcentcolon=P^{\prime}(\rho)/\rho$ and $q_{+}\vcentcolon=q(\rho_{+})$, and
$f^{0}\vcentcolon=$ $-\phi\textnormal{{div}}\psi+(\rho_{+}-\tilde{\rho})\textnormal{{div}}\psi-\psi\cdot\nabla\tilde{\rho}-\phi\textnormal{{div}}\tilde{u},$ (3.8c)
$f^{i}\vcentcolon=$ $-\frac{\phi+\tilde{\rho}-\rho_{+}}{\rho_{+}(\phi+\tilde{\rho})}\big{\{}\mu\Delta\psi^{i}+(\mu+\lambda)\partial_{x_{i}}\textnormal{{div}}\psi\big{\}}+\tilde{f}^{i},$ (3.8d)
@@ 3 A-Priori Estimates
+$+\{q_{+}-q(\rho)\}\partial_{x_{i}}\phi-\frac{P^{\prime}(\rho_{+})-P^{\prime}(\rho)}{\phi+\tilde{\rho}}\partial_{x_{i}}\tilde{\rho},$ (3.8e)
for $i=1,2,3$. In addition, we denote that $\mathcal{L}\equiv\hat{e}_{1}\mathcal{L}^{1}+\hat{e}_{2}\mathcal{L}^{2}+\hat{e}_{3}\mathcal{L}^{3}$ as the vector-valued operator, and $f\equiv\hat{e}_{1}f^{1}+\hat{e}_{2}f^{2}+\hat{e}_{3}f^{3}$.
Moreover, we define the energy norm:
$N^{2}(T)\vcentcolon=$ $\sup\limits_{0\leq t\leq T}\big{\{}\|(\phi,\psi)(t,\cdot)\|_{3}^{2}+\|\partial_{t}\phi(t,\cdot)\|_{2}^{2}+\|\partial_{t}\psi(t,\cdot)\|_{1}^{2}\big{\}}$
$+\int_{0}^{T}\!\big{\{}\|\nabla\phi\|_{2}^{2}+\|\partial_{t}\phi\|_{3}^{2}+\|\nabla\psi\|_{3}^{2}+\|\partial_{t}\psi\|_{2}^{2}\big{\}}\,{\mathrm{d}}t.$ (3.9)
To perform the continuity argument using a-priori estimates, we need the existence and uniqueness of a local-in-time solution. This is stated in the following lemma.
@@
Suppose the initial data $(\rho_{0},u_{0})$ satisfies $\|(\phi_{0},\psi_{0})\|_{3}<\infty$. Then there exists $T_{\ast}>0$, depending only on $\|(\phi_{0},\psi_{0})\|_{3}$, such that the initial-boundary value problem (3.8) and (1.2)–(1.4) has a unique solution $(\rho,u)$ satisfying $N(T)<\infty$ for $T\in[0,T_{\ast})$.
The proof of Lemma 3.1 is obtained using the same procedure of [14]. It is a standard argument for the local-in-time well-poseness theorem, and we omit the detail in the present paper.
Given initial norm $N(0)>0$. By continuity of $T\mapsto N(T)$, there exists a time $T_{0}>0$ such that,
$N(T)\leq 2N(0)\qquad\text{for all }\ T\in[0,T_{0}].$ (3.10)
Then by the Sobolev embedding theorem $H^{2}(\Omega)\xhookrightarrow{}\mathcal{C}^{0}(\Omega)$, there exists a positive constant $\varepsilon=\varepsilon(\rho_{+},u_{b},\gamma,K,\mu,\lambda)$ such that if $|u_{b}|+N(0)\leq\varepsilon$ then
$\frac{\rho_{+}}{2}\leq\rho(t,x)\leq\frac{3}{2}\rho_{+}\qquad\text{for }\ (t,x)\in[0,T_{0}]\times\Omega.$ (3.11)
Moreover, it also follows by the Sobolev embedding theorem that there exists a generic constant $C>0$ such that for all $T>0$,
@@
The main aim of this section is derive the following theorem:
@@
Let $(\tilde{\rho},\tilde{u})$ be the stationary solution to (1.5). Suppose $(\rho,u)$ is the solution to the problem (1.1)–(1.4) with maximum time of existence $T_{\ast}>0$. There exist constants $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$, $\varepsilon=\varepsilon(\rho_{+},\gamma,K,\mu,\lambda)>0$, and $T_{0}=T_{0}(\varepsilon)\in(0,T_{\ast}]$ which are independent of $u_{b}$ and $(\rho_{0},u_{0})$ such that, if $|u_{b}|+N(0)\leq\varepsilon$, then
$N(T_{0})\leq CN(0).$
3.1 $L^{2}$ Energy Estimate
For $\zeta$, $\xi\geq 0$, we define the potential energy density $H(\zeta,\xi)$ to be
$H(\zeta,\xi)\!\vcentcolon=\!\zeta\!\int_{\xi}^{\zeta}\!\dfrac{P(z)-P(\xi)}{z^{2}}\,{\mathrm{d}}z\!=\!\left\{\begin{aligned} &\dfrac{K}{\gamma-1}\big{\{}\zeta^{\gamma}-\xi^{\gamma}-\gamma\xi^{\gamma-1}(\zeta-\xi)\big{\}}&&\text{if }\ \gamma>1,\\ &K\xi\big{(}1-\dfrac{\zeta}{\xi}+\dfrac{\zeta}{\xi}\ln\dfrac{\zeta}{\xi}\big{)}&&\text{if }\ \gamma=1.\end{aligned}\right.$ (3.13)
It can be verified that $H(\zeta,\xi)$ is a positive convex function achieving global minimum at $\{\zeta=\xi\}$ with $H(\zeta,\xi)=0$ if and only if $\zeta=\xi$. Moreover $H(\zeta,\xi)$ satisfies
$\zeta\partial_{\zeta}H(\zeta,\xi)=H(\zeta,\xi)+P(\zeta)-P(\xi),\qquad\xi\partial_{\xi}H(\zeta,\xi)=-P^{\prime}(\xi)(\zeta-\xi),$ (3.14a)
@@ 3.1 $L^{2}$ Energy Estimate
With $H(\cdot,\cdot)$, we define the $L^{2}$ relative energy density $\mathcal{E}$ to be:
:$\mathcal{E}\vcentcolon=\dfrac{1}{2}\rho|u-\tilde{u}|^{2}+H(\rho,\tilde{\rho}),\qquad\mathcal{E}_{0}\equiv\mathcal{E}|_{t=0}.$ (3.15)
Using Taylor’s Theorem, we obtain the following proposition,
@@
for some constant $C=C(\rho^{\ast},\rho_{\ast},\gamma,K)>0$.
@@
for all $T\in[0,T_{0}]$, where $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ is some constant independent of $u_{b}$.
@@
$=$ $(1-\gamma)H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}-\textnormal{{div}}\big{(}uH(\rho,\tilde{\rho})\big{)}-\dfrac{\phi}{\tilde{\rho}}(\psi\cdot\nabla)P(\tilde{\rho}).$
Integrating the above in $x\in\Omega$, and using the boundary condition (1.2a), we obtain
@@
:$\mathcal{F}\vcentcolon=\rho u\dfrac{|\psi|^{2}}{2}+\psi\big{(}P(\rho)-P(\tilde{\rho})\big{)}-\mu(\nabla\psi)^{\top}\cdot\psi-(\mu+\lambda)\psi\textnormal{{div}}\psi.$
Integrating the above in $x\in\Omega$, and using the boundary condition (1.2), we have
@@
Next, by the steady state momentum equation 1.5b, it follows that
@@
Combining (3.1) and (3.1), then using definition 3.15, we obtain
@@
$=$ $(1-\gamma)\int_{\Omega}\!\!H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x-\int_{\Omega}\!\!\big{\{}\rho\psi^{\top}\cdot\nabla\tilde{u}\cdot\psi+\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\big{\}}\,{\mathrm{d}}x$ (3.18)
Since $\tilde{u}=\frac{x}{r}\tilde{U}(r)$, for $r=|x|$, it follows that $\nabla\tilde{u}$ is a symmetric matrix and
$\nabla\tilde{u}=(\nabla\tilde{u})_{+}+(\nabla\tilde{u})_{-}\quad\text{and}\quad\textnormal{{div}}\tilde{u}=\partial_{r}\tilde{U}(r)+(n-1)\dfrac{\tilde{U}(r)}{r},$
@@
$(\nabla\tilde{u})_{+}\vcentcolon=\big{(}\partial_{r}\tilde{U}(r)-\dfrac{\tilde{U}(r)}{r}\big{)}\dfrac{x\otimes x}{r^{2}},\qquad(\nabla\tilde{u})_{-}\vcentcolon=\dfrac{\tilde{U}(r)}{r}\mathbb{I}_{3}.$
By Lemma 1.1, one can show that
@@
Moreover, using the above and Lemma 1.1, it also follows that for all $V\in\mathbb{R}^{3}$,
@@
Using this in (3.1), we obtain
@@
$=$ $-\int_{\Omega}\!\!\rho\psi^{\top}\cdot(\nabla\tilde{u})_{-}\cdot\psi\,{\mathrm{d}}x-\int_{\Omega}\!\!\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\,{\mathrm{d}}x$ (3.20)
By Lemma 1.1, (3.11), Propositions 3.3 and A.1, we have
@@
where we choose $|u_{b}|\leq\mu/(36\rho_{+})$. In addition, by Lemma 1.1,
$\textnormal{{div}}\tilde{u}=\tilde{\rho}(1)|u_{b}|\dfrac{\partial_{r}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}\geq\frac{|u_{b}|^{3}}{Cr^{7}},\quad|L\tilde{u}|=\mu\tilde{\rho}(1)|u_{b}|\cdot\Big{|}\dfrac{\partial_{r}^{2}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}-2\dfrac{\partial_{r}\tilde{\rho}}{r^{3}\tilde{\rho}^{2}}-2\dfrac{|\partial_{r}\tilde{\rho}|^{2}}{r^{2}\tilde{\rho}^{3}}\Big{|}\leq\frac{C|u_{b}|^{3}}{r^{8}}.$
Using the above, and Proposition 3.3, we obtain that
$(\gamma-1)\int_{\Omega}H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x\geq\frac{|u_{b}|^{3}}{C_{1}}\int_{\Omega}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x,$ (3.22)
for some constant $C_{1}=C_{1}(\rho_{+},\mu,\gamma,K)>0$. Moreover, by Lemma 1.1, Cauchy-Schwarz’s inequality, and Proposition A.1, we get
@@
where we choose $|u_{b}|$ small enough such that $C|u_{b}|^{3}\leq 1/C_{1}$. Substituting (3.1)–(3.23) into (3.1), we obtain
@@
Integrating the above in time, we obtain the desired estimate. $\square$

### v2

@@ 2.1 Reformulation and energy norm $N(T)$
[…]
Then the system (1.1)-(1.3) is rewritten to that of the unknown functions $(\phi,\psi)$ as
$\psi|_{\partial\Omega}=$ 0 $0\hskip 18.49988pt\text{ for }\ t\geq$ 0, (2.1a)
$\mathcal{L}^{0}(\phi,\psi)\vcentcolon$= $D_{t}\phi+\rho_{+}\textnormal{{div}}\psi=f^{0},$ (2.1b)
$\mathcal{L}^{i}(\phi,\psi)\vcentcolon$= $\partial_{t}\psi^{i}-\frac{\mu}{\rho_{+}}\Delta\psi^{i}-\frac{\mu+\lambda}{\rho_{+}}\partial_{x_{i}}\textnormal{{div}}\psi+q_{+}\partial_{x_{i}}\phi=f^{i},$ (2.1c)
[…]
$f^{0}\vcentcolon$= $-\phi\textnormal{{div}}\psi+(\rho_{+}-\tilde{\rho})\textnormal{{div}}\psi-\psi\cdot\nabla\tilde{\rho}-\phi\textnormal{{div}}\tilde{u}$, (2.1d)
$f^{i}\vcentcolon$= $-\frac{\phi+\tilde{\rho}-\rho_{+}}{\rho_{+}(\phi+\tilde{\rho})}\big\{\mu\Delta\psi^{i}+(\mu+\lambda)\partial_{x_{i}}\textnormal{{div}}\psi\big\}+\tilde{f}^{i},$ (2.1e)
@@
$+\{q_{+}-q(\rho)\}\partial_{x_{i}}\phi-\frac{P^{\prime}(\rho_{+})-P^{\prime}(\rho)}{\phi+\tilde{\rho}}\partial_{x_{i}}\tilde{\rho}$. (2.1f)
In addition, we define $\mathcal{L}:=\hat{e}_{1}\mathcal{L}^{1}+\hat{e}_{2}\mathcal{L}^{2}+\hat{e}_{3}\mathcal{L}^{3}$ as the vector-valued operator and $f:=\hat{e}_{1}f^{1}+\hat{e}_{2}f^{2}+\hat{e}_{3}f^{3}$.
Moreover, for given initial data $(\rho_{0},u_{0})$ and time $T>0$, we define the solution space $X(T)$ as
$X(T)\!\vcentcolon=\!\bigg\{(\phi,\psi)\!\in\!L^{\infty}\big(0,T;H^{3}(\Omega)\big)\,\bigg|\,\begin{aligned} &\phi\!\in\!H^{1}\big(0,T;H^{2}(\Omega)\big),\,\psi\!\in\!H^{1}\big(0,T;H^{1}(\Omega)\big),\\ &\psi|_{\partial\Omega}\!=\!0,\,(\phi,\psi)|_{t=0}\!=\!(\phi_{0},\psi_{0}),\,N(T)<\infty.\end{aligned}\bigg\},$
where the energy norm $N(T)$ is defined by
$N(T)\vcentcolon$= $\sup\limits_{0\leq t\leq T}\Big\{\|(\phi,\psi)(t,\cdot)\|_{3}+\|\partial_{t}\phi(t,\cdot)\|_{2}+\|\partial_{t}\psi(t,\cdot)\|_{1}\Big\}$
$+\bigg\{\int_{0}^{T}\!\big\{\|\nabla\phi\|_{2}^{2}+\|\partial_{t}\phi\|_{3}^{2}+\|\nabla\psi\|_{3}^{2}+\|\partial_{t}\psi\|_{2}^{2}\big\}\,{\mathrm{d}}t\bigg\}^{\frac{1}{2}}.$ (2.2)
To obtain the global-in-time well-posedness and time-asymptotic stability, we need the existence and uniqueness of a local-in-time solution. This is stated as follows.
@@
[…]
The proof of Lemma 2.1 follows the same procedure in [18], and so we omit the details. We mention that for the Cauchy problem, an analogous method is employed in [17] to obtain the local-in-time well-posedness(See Theorem 5.2 of [17]).
For initial data with $N_{0}\equiv N(0)<\infty$, there exists a time $T=T(N_{0})>0$ such that
$N(t)\leq 2N(0)\qquad\text{for }\ t\in[0,T]$ (2.3)
by the continuity of $t\mapsto N(t)$. By the Sobolev embedding theorem $H^{2}(\Omega)\xhookrightarrow{}\mathcal{C}^{0}(\Omega)$, there exists a constant $\varepsilon=\varepsilon(\rho_{+},u_{b},\gamma,K,\mu,\lambda)>0$ such that if $|u_{b}|+N(0)\leq\varepsilon$ then
$\frac{\rho_{+}}{2}\leq\rho(t,x)\leq\frac{3}{2}\rho_{+}\qquad\text{for }\ (t,x)\in[0,T]\times\Omega.$ (2.4)
Moreover, it also follows by the Sobolev embedding theorem that there exists a constant $C=C(\rho_{+},u_{b},\gamma,K,\mu,\lambda)>0$ such that,
@@
The main aim of Section 2 is to derive the following a-priori estimate:
@@
There exists two positive constants $\varepsilon_{0}=\varepsilon_{0}(\rho_{+},\gamma,K,\mu,\lambda)>0$ and $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$, which are independent of $|u_{b}|$, such that the following statement holds: if $(\rho,u)$ solves the problem (1.1)–(1.4) in some time interval $t\in[0,T]$ and its energy norm and boundary data satisfy $|u_{b}|+N(T)\leq\varepsilon_{0}$, then $N(T)\leq CN(0).$
If Lemma 2.1 and Theorem 2.2 are obtained, then the global existence of $(\rho,u)$ is proved by the standard continuity argument. For the detail, we refer readers to Section 3 of [19]. Moreover, the asymptotic convergence (1.8) is derived by the same argument presented in Theorem 7.1 of [17].
2.2 $L^{2}$-estimate
To show the a-priori estimate, we first derive the $L^{2}$-estimate of the perturbation $(\phi,\psi)$. To this end, we employ a potential energy density $H(\zeta,\xi)$, which is defined by
$H(\zeta,\xi)\!\vcentcolon=\!\zeta\!\int_{\xi}^{\zeta}\!\dfrac{P(z)-P(\xi)}{z^{2}}\,{\mathrm{d}}z\!=\!\left\{\begin{aligned} &\dfrac{K}{\gamma-1}\big\{\zeta^{\gamma}-\xi^{\gamma}-\gamma\xi^{\gamma-1}(\zeta-\xi)\big\}&&\text{if }\ \gamma>1,\\ &K\xi\big(1-\dfrac{\zeta}{\xi}+\dfrac{\zeta}{\xi}\ln\dfrac{\zeta}{\xi}\big)&&\text{if }\ \gamma=1,\end{aligned}\right$. (2.6)
where $\zeta$, $\xi\geq 0$. It is holds that $H(\zeta,\xi)$ is a positive convex function achieving global minimum at $\{\zeta=\xi\}$ with $H(\zeta,\xi)=0$ if and only if $\zeta=\xi$. Moreover $H(\zeta,\xi)$ satisfies
$\zeta\partial_{\zeta}H(\zeta,\xi)=H(\zeta,\xi)+P(\zeta)-P(\xi),\hskip 18.49988pt\xi\partial_{\xi}H(\zeta,\xi)=-P^{\prime}(\xi)(\zeta-\xi),$ (2.7a)
@@ 2.2 $L^{2}$-estimate
A $L^{2}$–relative energy density $\mathcal{E}$ is defined by
$\mathcal{E}\vcentcolon=\dfrac{1}{2}\rho|u-\tilde{u}|^{2}+H(\rho,\tilde{\rho}),\qquad\mathcal{E}_{0}:=\mathcal{E}|_{t=0}.$ (2.8)
By Taylor’s theorem, we have the following proposition.
@@ A $L^{2}$–relative energy density $\mathcal{E}$ is defined by
for a certain constant $C=C(\rho^{\ast},\rho_{\ast},\gamma,K)>0$.
@@
for $T\in[0,T_{0}]$, where $T_{0}$ is given in (2.3)–(2.4), and $C=C(\rho_{+},\gamma,K,\mu,\lambda)>0$ is a certain constant independent of $u_{b}$ and $(\rho_{0},u_{0})$.
@@
[…]
Integrating the above equality over $\Omega$ with using the boundary condition (1.2a), we obtain
@@
$\mathcal{F}\vcentcolon=\rho u\dfrac{|\psi|^{2}}{2}+\psi\big(P(\rho)-P(\tilde{\rho})\big)-\mu(\nabla\psi)^{\top}\cdot\psi-(\mu+\lambda)\psi\textnormal{{div}}\psi.$
Integrating the above equality over $\Omega$ with using the boundary condition (1.2), we have
@@
The integrands of the left hand of (2.2) is rewritten as
@@
The above equality is obtained from the stationary momentum equation (1.5b). Adding equation (2.2) to (2.2) with using definition 2.8, we obtain that
@@
$=$ $(1-\gamma)\int_{\Omega}\!\!H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x-\int_{\Omega}\!\!\big\{\rho\psi^{\top}\cdot\nabla\tilde{u}\cdot\psi+\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\big\}\,{\mathrm{d}}x.$ (2.11)
The right hand side of (2.2) is handled as follows. As $\tilde{u}=\frac{x}{r}\tilde{U}(r)$ for $r=|x|$, $\nabla\tilde{u}$ is a symmetric matrix and satisfies
$\nabla\tilde{u}=(\nabla\tilde{u})_{+}+(\nabla\tilde{u})_{-}\qquad\text{and}\qquad\textnormal{{div}}\tilde{u}=\partial_{r}\tilde{U}(r)+(n-1)\dfrac{\tilde{U}(r)}{r},$
@@
$(\nabla\tilde{u})_{+}\vcentcolon=\big(\partial_{r}\tilde{U}(r)-\dfrac{\tilde{U}(r)}{r}\big)\dfrac{x\otimes x}{r^{2}},\hskip 18.49988pt(\nabla\tilde{u})_{-}\vcentcolon=\dfrac{\tilde{U}(r)}{r}\mathbb{I}_{3}.$
By Lemma 1.1, we have the inequality
@@
Hence using the above inequalities and Lemma 1.1, we have that for an arbitrary vector $V\in\mathbb{R}^{3}$,
@@
Substituting this inequality in (2.2), we obtain that
@@
$\leq$ $-\int_{\Omega}\!\!\rho\psi^{\top}\cdot(\nabla\tilde{u})_{-}\cdot\psi\,{\mathrm{d}}x-\int_{\Omega}\!\!\phi\psi\cdot\dfrac{L\tilde{u}}{\tilde{\rho}}\,{\mathrm{d}}x.$ (2.13)
We estimate the right hand side of (2.2) term by term. By Lemma 1.1, (2.4), Propositions 2.3 and A.1, the first term is estimated as
@@
where we have chosen as $|u_{b}|\leq\mu/(36\rho_{+})$. In addition, It holds by Lemma 1.1 that
$\textnormal{{div}}\tilde{u}=\tilde{\rho}(1)|u_{b}|\dfrac{\partial_{r}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}\geq\frac{|u_{b}|^{3}}{Cr^{7}},\qquad|L\tilde{u}|=\mu\tilde{\rho}(1)|u_{b}|\cdot\Big|\dfrac{\partial_{r}^{2}\tilde{\rho}}{r^{2}\tilde{\rho}^{2}}-2\dfrac{\partial_{r}\tilde{\rho}}{r^{3}\tilde{\rho}^{2}}-2\dfrac{|\partial_{r}\tilde{\rho}|^{2}}{r^{2}\tilde{\rho}^{3}}\Big|\leq\frac{C|u_{b}|^{3}}{r^{8}}.$
By using the above inequality and Proposition 2.3, we obtain that
$(\gamma-1)\int_{\Omega}H(\rho,\tilde{\rho})\textnormal{{div}}\tilde{u}\,{\mathrm{d}}x\geq\frac{|u_{b}|^{3}}{C_{1}}\int_{\Omega}\dfrac{|\phi|^{2}}{|x|^{7}}\,{\mathrm{d}}x$ (2.15)
for a certain constant $C_{1}=C_{1}(\rho_{+},\mu,\gamma,K)>0$. Moreover, owing to Lemma 1.1, Cauchy-Schwarz’s inequality and Proposition A.1, we get
@@
where we have chosen $|u_{b}|$ so small that $C|u_{b}|^{3}\leq 1/C_{1}$. Substituting (2.2)–(2.16) in (2.2), we obtain
@@
Integrating the above inequality in time, we obtain the desired estimate. $\square$